OpenStudy (anonymous):

To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended 5 sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. Write the solution as an ordered pair.

OpenStudy (dumbcow):

membership fee is paid initially then cost is per session m+5f = 25 10f = 40 find f first f = 10/40 f=4 substitute this into first equation to find m m + 5(4) =25 m = 25 -20 m = 5 (5,4)